# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### 16 Mar 2026 v4

@@ Abstract
Within the prevailing theoretical paradigm, quantum critical fluctuations are assumed to be overdamped by fast electronic continua [3, 4]. This framework, formalized in Hertz–Millis theory, predicts a universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics. While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are ubiquitously observed [5, 6, 7, 8, 9].
@@
In this work we show that infrared collective dynamics is generically controlled by the TDOS of correlated internal decay channels. By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [15, 16, 17], we derive an exact spectral representation of the collective susceptibility in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
@@
We demonstrate that the infrared scaling of the relaxation-rate density of states, $\rho(\lambda)\sim\lambda^{\alpha}$, naturally classifies universality classes of dynamical critical behavior. Within the relaxation-spectrum framework, this scaling directly determines the low-frequency structure of the retarded susceptibility. While the conventional Hertz–Millis kernel $\mathrm{Im}\,\chi^{-1}\propto|\omega|$ corresponds to the case $\alpha=1$, arising microscopically from particle–hole polarization of a Fermi liquid, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic kernel proportional to $\mathrm{sgn}(\omega)$. In the time domain this corresponds to a universal long–memory response $K(t)\sim 1/t$. The latter regime reflects a fundamentally different physical mechanism: rather than overdamped order-parameter fluctuations generated by a fermionic bath, the infrared dynamics is governed by the collective organization of the relaxation spectrum itself. This memory-dominated regime therefore defines a distinct universality class controlled by the structure of the TDOS.
@@
Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [18, 19, 20]. The resulting long-memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a route to strong superconductivity without invoking finely tuned bosonic mediators.
@@
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 21, 22]. Charge, spin, and pairing channels reorganize within a common slow-mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a shared infrared spectral continuum.
@@
Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified spectral perspective on collective behavior across strongly correlated systems [5, 7, 10].
@@
The remainder of this paper is organized as follows. In Section II we derive the exact TDOS representation of the collective susceptibility and classify infrared universality classes. In Sections III and IV we analyze the resulting memory–dominated pairing kernel and its implications for superconducting instabilities. In Section V we discuss experimental consequences including superconducting domes, Uemura scaling, and dynamical anomalies. We conclude in Section VI with a summary and outlook.
@@
Figure 1: Dynamical universality map of correlated quantum matter. (a) Infrared time–scale density of states $\rho(\lambda)$ defining distinct dynamical universality classes. Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, while a flat TDOS defines memory–dominated criticality. (b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory. (c) Physical consequences of memory–dominated criticality, including algebraic pairing enhancement, superconducting domes, Uemura scaling, strange metal transport, and long–time noise correlations.
@@
define the fundamental decay channels of the interacting system. For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as
$\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0.$ (3)
Infrared dynamics is therefore organized by the distribution of these decay scales rather than by Hamiltonian excitation energies.
@@
To compute collective response functions, we employ a functional representation of the dissipative dynamics. The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework [23, 24]. Upon projection onto slow collective operators and in the regime where decoherence suppresses quantum interference between forward and backward contours, the theory reduces to a classical MSRJD functional integral describing the infrared stochastic dynamics.
@@
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda},$ (7)
@@
Equation (16) provides an exact infrared response fully controlled by the relaxation-rate density. The slow-mode reservoir thus appears as a memory-dominated fixed point, whose long-time kernel structure governs the dynamical amplification mechanism discussed below.
The explicit Gaussian integration and derivation of the memory kernel are presented in Appendix B.
@@
where the TDOS scaling directly implies the identification $s=\alpha$.
While this form resembles the conventional Hertz–Millis representation, its physical interpretation is different. In the present framework the exponent $s$ does not originate from fermionic damping of a single order–parameter mode but instead reflects the infrared organization of the entire relaxation spectrum.
@@
This reorganization of the collective spectrum qualitatively alters the structure of the Cooper channel. In conventional approaches superconductivity is typically attributed to the exchange of a specific bosonic “pairing glue”. In contrast, in the present framework superconductivity emerges from the dynamical amplification of pairing correlations generated by the slow–mode reservoir itself.
@@
where the $|\omega|$ term arises from Landau damping by the electronic continuum. This object describes the retarded susceptibility of an overdamped order-parameter fluctuation and should not be identified with the Cooper-channel polarization kernel itself.
In bosonic-mediator theories [27], the overdamped collective mode is then treated as an effective pairing mediator. After momentum integration, the resulting interaction kernel entering the Cooper channel is schematically of the form
$V(\omega)\sim\frac{1}{|\omega|},$ (26)
up to model-dependent form factors and ultraviolet regularization. If one further assumes that pairing can still be analyzed within a quasiparticle-based ladder or gap-equation framework, the linearized pairing equation takes the schematic form
@@
This benchmark should be distinguished sharply from the framework developed in the present work. Here we do not assume a well-defined quasiparticle description nor an effective pairing glue represented by a single bosonic propagator. Instead, after integrating out microscopic electronic degrees of freedom, the infrared Cooper sector is organized by the collective relaxation spectrum itself. The relevant object is not the retarded propagator of an overdamped order-parameter mode, but the polarization kernel generated by a continuum of relaxational pair modes, as derived below.
III.2 B. Collective relaxation spectrum and TDOS representation
Strongly correlated quantum matter, however, can develop a much richer infrared structure. Instead of a single overdamped bosonic excitation, the system may host a dense continuum of slow collective modes.
The collective dynamics can then be represented by the TDOS $\rho(\lambda)$, which measures the density of relaxation rates $\lambda$. Projecting the dynamics onto the slow collective sector yields the spectral representation
$\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega}.$ (31)
While the quantity above governs the retarded response of the collective sector, the object relevant for superconducting pairing is the Cooper-channel fluctuation kernel. After integrating out microscopic electronic degrees of freedom, pairing correlations are encoded in the two-point correlator of the collective pair field
$\langle\Delta(\omega)\Delta^{*}(\omega)\rangle.$ (32)
To understand its structure, consider a single relaxational pair mode with decay rate $\lambda$. At long times its dynamics is well approximated by a linear stochastic equation
$\dot{\Delta}_{\lambda}=-\lambda\,\Delta_{\lambda}+\xi_{\lambda},$ (33)
where $\xi_{\lambda}$ denotes Gaussian noise generated by microscopic degrees of freedom. The stationary two-point correlation function of such a mode is
$\langle\Delta_{\lambda}(t)\Delta_{\lambda}^{*}(0)\rangle\propto e^{-\lambda|t|}.$ (34)
Fourier transforming this correlation function gives the fluctuation spectrum of a relaxational mode
$S_{\lambda}(\omega)\propto\frac{1}{\lambda^{2}+\omega^{2}}.$ (35)
The Cooper-channel polarization kernel therefore arises by summing the contributions of all relaxational pair modes weighted by the TDOS,
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (36)
Equation (36) represents the dynamical polarization of the Cooper channel generated by the collective relaxation spectrum. Importantly, this object differs from the reactive part of the retarded susceptibility discussed above. Rather than describing the propagation of a single bosonic mediator, it measures the fluctuation spectrum of the entire slow-mode reservoir.
For a flat TDOS in the infrared,

### 30 Mar 2026 v5

@@ Abstract
Within the prevailing theoretical paradigm, quantum critical fluctuations are assumed to be overdamped by fast electronic continua [3, 4]. This framework, formalized in Hertz–Millis theory, predicts an approximately universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics. While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are widely observed [5, 6, 7, 8, 9, 10].
@@
In this work we show that infrared collective dynamics is generically controlled by the TDOS of correlated internal decay channels. Starting from the real–time Keldysh functional integral formulation of dissipative many-body dynamics [16, 17, 18], and passing to its classical stochastic infrared limit described by the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [19, 20, 21], we construct a dynamical coarse–graining theory in which the collective susceptibility admits an exact spectral representation in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
@@
We demonstrate that the infrared scaling of the relaxation-rate density of states, $\rho(\lambda)\sim\lambda^{\alpha}$, naturally classifies universality classes of dynamical critical behavior. Within the relaxation-spectrum framework, this scaling directly determines the low-frequency structure of the retarded susceptibility. While the conventional Hertz–Millis kernel $\mathrm{Im}\,\chi^{-1}\propto|\omega|$ corresponds to an effective TDOS scaling with $\alpha=1$, a flat TDOS ($\alpha=0$) instead produces a marginal nonanalytic response characterized by logarithmic frequency dependence and long-time memory. In the time domain this corresponds to a universal long–memory response $K(t)\sim 1/t$. The latter regime reflects a fundamentally different physical mechanism: rather than overdamped order-parameter fluctuations generated by a fermionic bath, the infrared dynamics is governed by the collective organization of the relaxation spectrum itself. This memory-dominated regime therefore defines a distinct universality class controlled by the structure of the TDOS.
@@
Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [22, 23, 24]. The resulting long-memory kernel enhances intrinsic electronic pairing tendencies, establishing a natural dynamical environment for strengthened pairing without relying on finely tuned bosonic mediators. Rather than arising from the formation of tightly bound Cooper pairs, superconductivity in this framework is understood as a collective dynamical phenomenon. A weak intrinsic pairing tendency generates fluctuating pairs, whose instability is driven by the accumulation of slow collective modes in the relaxation spectrum.
@@
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing tendencies, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 25, 26]. Charge, spin, and pairing channels reorganize within a common slow-mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a shared infrared spectral continuum.
@@
Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a natural dynamical environment for enhanced pairing and superconducting instabilities. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified spectral perspective on collective behavior across strongly correlated systems [5, 7, 11].
@@
The remainder of this paper is organized as follows. In Section II we derive the exact TDOS representation of the collective susceptibility and classify infrared universality classes. In Sections III and IV we analyze the resulting memory–dominated pairing kernel and its implications for superconducting instabilities. In Section V we discuss experimental consequences including superconducting domes, Uemura scaling, and dynamical anomalies. In Section VI we place these results in a broader conceptual framework, discussing their implications for dynamical criticality, Liouvillian descriptions of collective dynamics, and routes toward enhanced superconductivity, before concluding with a summary and outlook.
@@
Figure 1: Dynamical universality map of correlated quantum matter. (a) Infrared time–scale density of states $\rho(\lambda)$ defining distinct dynamical universality classes. Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, a flat TDOS defines memory–dominated criticality, and singular infrared spectra ($\alpha<0$) correspond to a dense accumulation of slow modes. (b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory, with $K(t)\sim 1/t$ for the marginal (flat TDOS) case and slower decay for IR-singular spectra. (c) Physical consequences of the TDOS-controlled dynamics. Memory–dominated criticality produces scale-free dynamics, long-time correlations, and logarithmic enhancement of the retarded pairing kernel, while additional infrared accumulation of slow modes ($\alpha<0$) leads to power-law singularities and algebraic enhancement of superconducting pairing.
@@
define the fundamental decay channels of the interacting system. For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as $\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0$. Infrared dynamics is therefore organized by the distribution of these decay scales rather than by Hamiltonian excitation energies.
@@
To compute collective response functions, we employ a functional representation of the dissipative dynamics. The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework. Upon projection onto slow collective operators and in the regime where decoherence suppresses quantum interference between forward and backward contours, the theory reduces to a classical MSRJD functional integral describing the infrared stochastic dynamics.
@@
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\xi_{\lambda},$ (6)
@@
Equation (15) provides an exact infrared response fully controlled by the relaxation-rate density. The slow-mode reservoir thus appears as a memory-dominated dynamical regime, whose long-time kernel structure governs the infrared dynamical behavior discussed below. The explicit Gaussian integration and derivation of the memory kernel are presented in Appendix B.
@@
where the TDOS scaling directly implies the identification $s=\alpha$. While this form resembles the conventional Hertz–Millis representation, its physical interpretation is different. In the present framework the exponent $s$ does not originate from fermionic damping of a single order–parameter mode but instead reflects the infrared organization of the entire relaxation spectrum.
@@
This reorganization of the collective spectrum qualitatively alters the structure of the Cooper channel. In conventional approaches superconductivity is typically attributed to the exchange of a specific bosonic “pairing glue”. In contrast, in the present framework superconductivity emerges from the infrared structure of pairing correlations reorganized by the slow-mode reservoir.
@@
where the $|\omega|$ term arises from Landau damping by the electronic continuum. This object describes the retarded susceptibility of an overdamped order-parameter fluctuation.
In bosonic-mediator theories [29], the overdamped collective mode is then treated as an effective pairing mediator. After momentum integration, the resulting interaction kernel entering the Cooper channel is schematically of the form $V(\omega)\sim 1/|\omega|$, up to model-dependent form factors and ultraviolet regularization. If one further assumes that pairing can still be analyzed within a quasiparticle-based ladder or gap-equation framework, the linearized pairing equation takes the schematic form
@@
This benchmark should be distinguished sharply from the framework developed in the present work. Here we do not assume a well-defined quasiparticle description nor an effective pairing glue represented by a single bosonic propagator. Instead, after integrating out microscopic electronic degrees of freedom, the infrared Cooper sector is organized by the collective relaxation spectrum itself. The relevant object is not the retarded propagator of an overdamped order-parameter mode, but by the retarded polarization kernel generated by a continuum of relaxational pair modes, as derived below.
III.2 B. Retarded Cooper kernel and ladder resummation in the relaxation-spectrum framework
In strongly correlated systems, the Cooper channel must be formulated directly in terms of collective response rather than quasiparticle propagators. After microscopic fermions are integrated out, the infrared dynamics of pairing is governed by the relaxation spectrum introduced in Sec. II.
The appropriate object controlling superconducting instability is the retarded response function of the collective pair field,
$L^{R}(\omega)=\frac{1}{\lambda-i\omega}\quad\text{(single relaxational mode)}.$ (29)
For a continuum of relaxational modes characterized by the TDOS $\rho(\lambda)$, the retarded Cooper-channel kernel is obtained by integrating over the spectrum,
$\Pi^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (30)
This expression has the same spectral structure as the retarded self-energy derived in Sec. II, reflecting that both quantities are governed by the same underlying relaxation spectrum, although they correspond to distinct physical responses.
This kernel represents the linear response of the collective pair field and plays the role of the elementary rung in the Cooper ladder. The full pair susceptibility is obtained by resumming repeated pair scattering processes,
$\chi_{\Delta}^{R}(\omega)=\Pi^{R}(\omega)+\Pi^{R}(\omega)\,g_{\rm pair}\,\chi_{\Delta}^{R}(\omega),$ (31)
which yields
$\chi_{\Delta}^{R}(\omega)=\frac{\Pi^{R}(\omega)}{1-g_{\rm pair}\Pi^{R}(\omega)},$ (32)
where $g_{\rm pair}$ denotes the intrinsic electronic interaction that seeds pairing.
Superconductivity corresponds to a pole of the retarded susceptibility, and the Cooper instability is determined by the Thouless criterion,
$1-g_{\rm pair}\,\mathrm{Re}\,\Pi^{R}(\omega)=0.$ (33)
Using Eq. (30), the real part of the kernel is
$\mathrm{Re}\,\Pi^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\;\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}}.$ (34)
For a power-law TDOS, $\rho(\lambda)\sim\lambda^{\alpha}$, the infrared behavior of the retarded kernel depends on the sign of $\alpha$. In the infrared-singular regime $\alpha<0$, one finds
$\mathrm{Re}\,\Pi^{R}(\omega)\sim|\omega|^{\alpha}\int_{0}^{\infty}dx\,\frac{x^{\alpha+1}}{x^{2}+1}=C_{\alpha}|\omega|^{\alpha},$ (35)
where $C_{\alpha}$ is a positive finite constant. This nonanalytic power-law behavior directly determines the structure of the superconducting instability.
The fluctuation spectrum of pairing correlations is encoded in the Keldysh component,
$L^{K}(\omega)=L^{R}(\omega)\,\Sigma^{K}(\omega)\,L^{A}(\omega),$ (36)
which captures infrared enhancement of fluctuations and long-time correlations. However, the superconducting instability itself is controlled by the retarded response through Eq. (33).
Thus, within the relaxation-spectrum framework, superconducting instability is governed by the structure of the retarded kernel, while fluctuation correlators describe the accompanying dynamical noise and non-Markovian correlations. A detailed derivation of the response and fluctuation kernels is presented in Appendix D.
III.3 C. Infrared scaling and renormalization of the Cooper channel
The superconducting instability is determined by the Thouless criterion derived in Sec. III B. The infrared behavior of pairing is therefore controlled by the low-frequency scaling of the retarded kernel $\Pi^{R}(\omega)$, which is in turn determined by the structure of the relaxation spectrum.
In the flat-TDOS regime ($\alpha=0$), the kernel takes the form
$\Pi^{R}(\omega)\simeq\rho_{0}\ln\frac{\Lambda}{|\omega|},$ (37)
so that coarse graining over a logarithmic frequency shell $\omega\in[\Lambda/b,\Lambda]$ produces
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega}{\omega}=g^{2}\ln b,$ (38)
leading to the marginal flow $dg/dl=Cg^{2}$. Thus, a flat TDOS yields a logarithmic infrared enhancement of pairing, analogous to marginal behavior in conventional Eliashberg theories.
When the TDOS develops an infrared divergence ($\alpha<0$), the retarded kernel acquires a power-law form,
$\Pi^{R}(\omega)\sim|\omega|^{\alpha}.$ (39)
Under the scale transformation $\omega\to\omega/b$, the kernel transforms as
$\Pi^{R}\to b^{-\alpha}\Pi^{R},$ (40)
implying a scale dependence of the effective coupling
$g^{\prime}=b^{-\alpha}g.$ (41)
The corresponding RG flow is therefore
$\frac{dg}{dl}=(-\alpha)\,g+Cg^{2},$ (42)
so that for $\alpha<0$ the interaction becomes relevant and grows algebraically under coarse graining.
The qualitative difference from the Eliashberg regime originates from the organization of the relaxation spectrum. In the marginal case, each logarithmic shell contributes independently, leading to only logarithmic enhancement. By contrast, IR-singular spectra correspond to an accumulation of slow modes toward vanishing relaxation rates, which reorganizes the Cooper channel into a relevant infrared interaction.
The corresponding RG flows are illustrated schematically in Fig. 2. The schematic curves correspond to the solutions of the RG equations $dg/dl=Cg^{2}$ (marginal case) and $dg/dl=(-\alpha)g$ (IR-singular case), yielding logarithmic and exponential growth of the effective pairing interaction, respectively. In the marginal regime associated with a flat TDOS ($\alpha=0$), the pairing interaction increases only logarithmically under coarse graining, reflecting the weak accumulation of infrared spectral weight. By contrast, for infrared-singular TDOS ($\alpha<0$), the coupling acquires a finite scaling dimension and grows exponentially with the RG scale, indicating that the Cooper channel becomes a relevant infrared instability. The figure thus provides a direct visualization of how the infrared structure of the relaxation spectrum controls the qualitative nature of pairing enhancement.
Superconductivity thus emerges as an instability of the retarded collective response controlled by the TDOS, with logarithmic enhancement for flat spectra and algebraic growth only in the presence of an IR-singular slow-mode density.
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. A flat relaxation-rate density of states (marginal TDOS, $\alpha=0$) produces logarithmic growth, $dg/dl\sim g^{2}$, corresponding to a marginal infrared enhancement of pairing (blue). This behavior is analogous to the logarithmic scaling familiar from conventional BCS/Eliashberg-type pairing theories. In contrast, IR-singular relaxation spectra with enhanced infrared weight ($\alpha<0$) generate a relevant contribution, $dg/dl=(-\alpha)g$, leading to algebraic growth of the pairing interaction (orange). .
IV IV. Pairing from Memory–Dominated Critical Dynamics
The previous section established the universal infrared structure of the Cooper channel generated by the relaxation spectrum. In particular, the slow–mode reservoir reorganizes the retarded Cooper kernel into a singular form, leading to marginal enhancement of pairing when the TDOS remains finite in the infrared, and to a relevant instability only when the TDOS develops an infrared divergence.
The remaining question is how this infrared amplification is tied to microscopic electronic pairing tendencies in correlated materials and how it determines the observable superconducting transition scale. In this section we connect the intrinsic electronic pairing seed to the memory–dominated Cooper kernel and show how their interplay produces an enhanced superconducting instability, which becomes algebraic only when the TDOS develops an infrared divergence.
In the present framework, superconductivity does not originate from the formation of tightly bound Cooper pairs at the microscopic level. Instead, a weak intrinsic pairing tendency is already present in the electronic system, giving rise to fluctuating pair fields. The key mechanism is the accumulation of slow collective modes in the relaxation spectrum, which enhances the lifetime and overlap of these fluctuations. Superconductivity thus emerges as a collective dynamical instability driven by the reorganization of relaxation modes, rather than by pair formation itself.
Equivalently, the superconducting transition should be understood not as a bound–state formation problem, but as a dynamical instability of the retarded Cooper kernel controlled by the infrared structure of the relaxation spectrum.
IV.1 A. Intrinsic electronic pairing and dynamical formulation
In strongly correlated materials, short-range electronic interactions naturally generate local singlet correlations, providing a microscopic seed for pairing. For example, in systems proximate to Mott insulating phases, superexchange processes arising from strong Coulomb repulsion favor spin-singlet formation [1, 32, 33]. More generally, however, the specific microscopic origin of pairing is not essential for the present analysis.
At low energies, interacting electrons may be described by a generic four–fermion interaction,
$H_{\rm int}=\sum_{k,k^{\prime},q}V(q)\,c^{\dagger}_{k+q\sigma}c^{\dagger}_{k^{\prime}-q\sigma^{\prime}}c_{k^{\prime}\sigma^{\prime}}c_{k\sigma},$ (43)
which contains multiple collective channels. Projecting this interaction onto the particle–particle sector defines an effective pairing vertex in the Cooper channel, which serves as the intrinsic pairing interaction $g_{\rm pair}$.
In conventional approaches these channels are analyzed through diagrammatic resummations, often emphasizing a specific bosonic mediator. In contrast, the present framework focuses on the dynamical organization of the collective spectrum itself.
While several ordering channels are in principle allowed, their infrared efficiency differs. Particle–hole instabilities such as density waves require specific momentum structure, restricting their available phase space. By contrast, superconductivity corresponds to condensation in the particle–particle channel at zero total momentum,
$\Delta=\langle c_{k\uparrow}c_{-k\downarrow}\rangle,$ (44)
and can therefore efficiently accumulate infrared spectral weight.
Within the dynamical framework developed in Sec. III, the pairing susceptibility assumes a universal tt–matrix structure, as given in Eq. (32). In this expression, $g_{\rm pair}$ represents the intrinsic electronic pairing seed, while $\Pi^{R}(\omega)$ denotes the retarded Cooper kernel generated by the relaxation spectrum after integrating out microscopic degrees of freedom.
A key point is that the present framework does not rely on a specific microscopic pairing mechanism. Instead, short-range electronic correlations provide a generic pairing seed, while the infrared dynamical spectrum controls how this tendency is amplified.
In weakly interacting systems the pairing instability remains marginal, leading to the exponentially small transition scale of conventional BCS theory. Near correlated criticality, however, the relaxation spectrum becomes dense near $\lambda\to 0$, generating an extensive reservoir of slow collective modes. This slow–mode reservoir reorganizes the infrared dynamics and strongly enhances the Cooper channel. Thus, critical dynamics does not create attraction, but amplifies a preexisting pairing tendency through infrared spectral reorganization.
IV.2 B. Fluctuation enhancement versus pairing instability
The analysis above reveals a crucial distinction between two closely related but physically different quantities: the fluctuation correlator of the pair field and the retarded response kernel that governs superconducting instability.
Within the Keldysh formulation, the fluctuation spectrum of pairing is encoded in the correlator
$\Pi^{K}(\omega)\equiv\langle\Delta(\omega)\Delta^{\dagger}(\omega)\rangle_{K},$ (45)
which, for a continuum of relaxational modes, takes the form
$\Pi^{K}(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm pair}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (46)
For a flat TDOS, $\rho_{\rm pair}(\lambda\to 0)=\rho_{0}$, this yields
$\Pi^{K}(\omega)\sim\frac{\rho_{0}}{|\omega|},$ (47)
indicating a strong infrared enhancement of pairing fluctuations. In the time domain, this corresponds to long-time correlations $K(t)\sim 1/t$ and scale-free noise.
However, the superconducting transition is not determined by the fluctuation correlator, but by the retarded response function through the Thouless criterion in Eq. (33). The retarded kernel exhibits a different infrared structure. For a flat TDOS, $\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln\frac{\Lambda}{|\omega|}$, leading to a logarithmic (marginal) enhancement and a BCS-like transition scale.
Thus, although pairing fluctuations are strongly enhanced, this enhancement alone does not produce an algebraic increase of the transition temperature. The transition is controlled by the infrared structure of the retarded kernel, not by the fluctuation amplitude itself.
A qualitatively different regime arises only when the TDOS develops an infrared divergence, $\rho_{\rm pair}(\lambda)\sim\lambda^{\alpha}(\alpha<0)$, for which $\mathrm{Re}\,\Pi^{R}(\omega)\sim|\omega|^{\alpha}$. In this case the retarded kernel itself becomes singular, and the pairing interaction acquires a relevant scaling dimension, leading to algebraic growth of the superconducting instability.
Physical implication. Strong infrared fluctuations, long-time memory, and $1/f$-type noise are therefore natural consequences of a flat TDOS, but they do not by themselves imply high transition temperatures. Enhanced superconductivity requires not only strong fluctuations, but an infrared-divergent accumulation of slow modes that modifies the retarded response itself.
This distinction clarifies that superconductivity in the present framework is controlled by the structure of the retarded kernel, while fluctuation correlators characterize the accompanying non-equilibrium and dynamical signatures.
IV.3 C. Transition scale from infrared spectral structure
The superconducting transition is determined by the Thouless criterion in Eq. (33), which shows that the transition temperature is controlled by the infrared behavior of the retarded Cooper-channel kernel.
At finite temperature, the imaginary-time direction becomes compact, $\tau\in[0,\beta]$ with $\beta=1/T$, and the system cannot probe arbitrarily long time scales. Equivalently, the lowest Matsubara frequency sets the infrared scale,
$|\omega|\sim T.$ (48)
Thus, thermal fluctuations provide an infrared cutoff, and the transition temperature is determined by evaluating $\Pi^{R}(\omega)$ at an effective scale.
Flat TDOS ($\alpha=0$): marginal regime. For a finite infrared TDOS,